OpenStudy (anonymous):

HELP FOR MEDAL? The surface area of a cube is given by the formula 6s^2, where s represents the length of each side of the cube. If the surface area of a cube is 160 square inches, what is the length of each side to the nearest tenth of an inch?

TheSmartOne (thesmartone):

\(\sf Surface~Area = 6s^2\) \(\sf 160 = 6s^2\)

TheSmartOne (thesmartone):

What is \(\sf\Large\frac{160}{6}=?\)

OpenStudy (anonymous):

rounded to the nearest tenth @TheSmartOne?

TheSmartOne (thesmartone):

Well we want the most exact answer at the end so we can round to at least 3 digits and then for the final answer round to the tenth.

OpenStudy (anonymous):

So \[\frac{ 160 }{ 6 } = 26.6\] Would that be right @TheSmartOne?

TheSmartOne (thesmartone):

Yes. And so now we have \(\sf 26.6 = s^2\) Now we have to take the squareroot of 26.6 And then we will get \(\sf\large\sqrt{26.6}=s\)

TheSmartOne (thesmartone):

Sorry, had lost internet connection and it has finally come back >.<